OpenStudy (anonymous):

square root of (3x+13) = square root of (20-4x)

OpenStudy (anonymous):

\[\sqrt{3x+13}=\sqrt{20-4x}\] Start by squaring both sides.

OpenStudy (lalaly):

square both sides 3x+13 = 20-4x 7x =7 x=1

OpenStudy (saifoo.khan):

\[\sqrt{3x+13} = \sqrt{20-4x}\]

OpenStudy (saifoo.khan):

LALA's right.

OpenStudy (anonymous):

by squaring both sides u get 3x+13=20-4x -x=7(by grouping like terms) x=-7

OpenStudy (lalaly):

groupin like terms you get 3x+4x = 20-13

OpenStudy (anonymous):

Me thinks you may want to recheck that addition Remainder :)

OpenStudy (anonymous):

uh yeah